# Post-glitch Recovery and the Neutron Star Structure: The Vela Pulsar

### v1

@@ IV.3 Correlations and Braking Index
To date, the Vela pulsar has exhibited 26 glitches, with 21 of them being with large magnitudes exceeding 10-8. Fig 11 presents these 21 events, revealing a pattern where glitch magnitudes exhibit an alternating increasing and decreasing trend: if glitch A has a lower magnitude, the successive glitch is observed with a higher magnitude, and vice versa. This trend is particularly evident in the last 10 glitches (or after MJD 50000).
@@
We also utilized the data of these 21 events to investigate potential correlations between the glitch amplitude and the inter-glitch time. Most correlations were not statistically significant except a correlation between glitch amplitude and subsequent inter-glitch time, shown in the Fig 12. Specifically, the amplitude of a glitch (e.g., glitch A) showed a positive correlation with the inter-glitch time following it (i.e., the time difference between glitch A+1 and glitch A). The correlation coefficients were: Pearson’s = 0.6582 (P–value = 0.0022), Spearman’s = 0.6266 (P–value = 0.0041), and Kendall’s = 0.4692 (P–value = 0.0051), indicating moderate to strong positive correlation. Such a tight correlation has only been found for PSR J0537–6910 (Ho et al., 2020).
@@
We measured the Vela pulsar’s braking index by estimating the frequency second derivative through linear fitting of frequency first derivative measurements just before glitch epochs, yielding a braking index value of 2.94 0.55. As a caution, it may be noted that our estimate assumes almost a full recovery with stable values of the first and second spin derivatives. This value is expected to become more constrained with future observations as this would increase glitch numbers or dataset coverage. It is close to the prediction of the model that invokes the magnetic dipole radiation in a plasma-filled magnetosphere (e.g. Ekşi et al. (2016)). Previous attempts to estimate the braking index for the Vela pulsar by different studies yielded qualitatively different results: a value of 1.4 0.2 (Lyne et al., 2010), 1.7 0.2 (Espinoza et al., 2017) and 2.81 0.12 (Akbal et al., 2017), respectively. The values in each case were different potentially attributed to differences in techniques used to estimate spin parameters and their derivatives and assumptions.
V Conclusion and Future Prospects
@@
The investigation of post-glitch recovery offers valuable insights into neutron star internal structure, which is not fully understood yet, and illuminates the underlying mechanisms driving these sudden spin-up events. Focusing on the Vela pulsar’s last four glitches, spanning from MJD 57651 to 60700 ( 100 months of observations), we employed the vortex creep model to extract key neutron star parameters. Bayesian analysis demonstrated that the recovery of the Vela glitches is well described by a hybrid model combining exponential and linear relaxations. Additionally, Bayesian inference closely predicted the time to the next glitch for all the glitches and suggested that the upcoming glitch is likely to occur around MJD 61377.7 (December 3, 2026), with a 68% credible interval spanning from MJD 61249 to 61506. The vortex creep model provided an excellent fit for the observed recoveries.
Furthermore, we analyzed the vortex residuals (the difference between the vortex creep model predictions and the measured spin-down rate) within the vortex bending model framework, conducting the first study of this kind for the Vela pulsar. Our analysis confirms damped quasi-periodic oscillations and yields precise determinations of the oscillatory parameters. Intriguingly, we also see an unexplained feature in the vortex residuals just preceding the subsequent glitch. These findings provide a foundational understanding of vortex residuals and motivate further enhancements in future both in theory and in similar follow-up investigations. Moreover, the measured parameters can improve the precision of hydrodynamical simulations of superfluid vortices, allowing for more accurate modeling and deeper insights into the neutron star’s dynamics and structural properties.
We explored several correlations and our analysis revealed a positive correlation between the glitch amplitude and the subsequent inter-glitch time, suggesting a potential link between the magnitude of glitches and the time elapsed before the next occurrence. Furthermore, we estimated the pulsar’s braking index as 2.94 0.55. As more glitches are observed, this estimate is likely to become more precise, yielding a more refined understanding of the pulsar’s rotational dynamics and offering deeper insights into its internal structure, magnetospheric conditions and evolutionary history.
@@
Future research includes investigating the post-glitch recovery and internal structure of other notable glitching pulsars, such as the Crab and PSR J1740-3015. To streamline the process, we are developing automated rotational parameter estimation scripts that perform the timing analysis quickly, enabling rapid model investigation and facilitating timely results. Additionally, the measured recovery parameters provide a unique approach to constrain the neutron star equation of state, which will be investigated in a subsequent work in progress. Furthermore, determining the internal and surface temperatures of neutron stars and exploring correlations with glitch observables could yield valuable insights. Lastly, constraining internal magnetic field configurations will also be a key area of investigation, ultimately advancing our understanding of neutron star physics and glitch mechanisms.
@@ Acknowledgements
EG is supported by the National Natural Science Foundation of China (NSFC) programme 11988101 under the foreign talents grant QN2023061004L. BCJ acknowledges the support from Raja Ramanna Chair fellowship of the Department of Atomic Energy, Government of India (RRC – Track I Grant 3/3401 Atomic Energy Research 00 004 Research and Development 27 02 31 1002//2/2023/RRC/R&D-II/13886). BCJ also acknowledges support from the Department of Atomic Energy Government of India, under project number 12-R&D-TFR-5.02-0700. PA acknowledges the support from SERB-DST, Govt. of India, via project code CRG/2022/009359.
@@
Data and codes used for this research will be provided on reasonable requests.
@@ Appendix A Results of Bayesian Analysis
Here, we present the outcomes of the Bayesian model comparison and the posteriors for the parameters estimated for each post-glitch recovery using the most preferred model. The results of the Bayesian Model comparison are listed in Tables 3 and 4. The posterior distributions are plotted with 68% and 95% credible intervals for all the cases given in Figs 13, 14, 15, 16, 17 and 18.
@@
| 57734 | 0 | 1350.0 | –3365.7 |
| Peak 1 Shifted | 0 | 730.9 | –3377.2 |
| 59417 | 0 | 4514.4 | –13296.9 |
| Peak 1 Shifted | 0 | –10572.9 | –13302.8 |
| Peak 2 Shifted | 0 | –3800.1 | –13301.3 |

### v2

@@ 4.3 Correlations and Braking Index
To date, 26 glitches have been reported in the Vela pulsar333http://www.jb.man.ac.uk/pulsar/glitches.html (Espinoza2011; Basu2022), with 21 of them having large magnitudes (¿ 10-7). Fig. 11 presents these 26 events. A pattern is observed among the large glitches, where their magnitudes exhibit an alternating trend: a glitch with a lower magnitude is typically followed by one with a higher magnitude, and vice versa. This alternating behaviour is particularly evident in the last 10 glitches (i.e., after MJD 50000), as highlighted by the red boundary in Fig. 11.
@@
We used the reported data of 26 glitch events to investigate potential correlations between the glitch magnitude and the inter-glitch time. A significant correlation was observed between the glitch magnitude and subsequent inter-glitch time specifically for events with large magnitudes, shown in Fig. 12. The correlation coefficients were: Pearson’s = 0.6470 (P–value = 0.0020), Spearman’s = 0.5852 (P–value = 0.0067), and Kendall’s = 0.4274 (P–value = 0.0086), indicating moderate to strong positive correlation. Such a correlation has only been found for PSR J0537–6910 (ho20).
This correlation between glitch magnitude and the subsequent interglitch time does not persist when small glitches are included, as illustrated in Fig. 13. The reality of reported small glitch events remains uncertain; some may be consistent with timing noise (Espinoza2011; Grover2024), although it is unlikely that all small glitches are timing noise, especially those accompanied by exponential recoveries (Espinoza2021). Therefore, the interpretation of this correlation should be approached with extreme caution. If the features and correlations shown in Figs. 11 and 12 persist in including all real glitches; this would have significant implications for glitch forecasting models, such as those proposed by Antonelli2022.
@@
We measured the Vela pulsar’s braking index by estimating the frequency second derivative through linear fitting of frequency first derivative measurements just before glitch epochs, yielding a braking index value of 2.94 0.55. As a caution, it may be noted that our estimate assumes almost a full recovery with stable values of the first and second spin derivatives. This value is expected to become more constrained with future observations as this would increase glitch numbers or dataset coverage. It is close to the prediction of the model that invokes the magnetic dipole radiation in a plasma-filled magnetosphere (e.g. eksi16). Previous attempts to estimate the braking index for the Vela pulsar by different studies yielded qualitatively different results: a value of 1.4 0.2 (Lyne1996), 1.7 0.2 (Espinoza2017) and 2.81 0.12 (Akbal2017), respectively. The values in each case were different, potentially attributed to differences in techniques used to estimate spin parameters and their derivatives and assumptions.
5 Conclusion and Future Prospects
@@
Focusing on the Vela pulsar’s last four glitches, spanning from MJD 57651 to 60700 ( 100 months of observations), we employed the vortex creep model to infer model parameters relevant to the post-glitch evolution. Bayesian analysis demonstrated that the recovery of the Vela glitches is well described by a hybrid model combining exponential and linear relaxations. Additionally, Bayesian inference closely predicted the observed time to the next glitch for all the glitches and suggested that the upcoming glitch is likely to occur around MJD 61377.7 (December 3, 2026), with a 68% credible interval spanning from MJD 61249 to 61506.
The vortex creep model provides a good description of the observed post-glitch recoveries. We further analysed the vortex residuals, defined as the difference between the model predictions and the measured spin-down rate, within the vortex bending framework. This analysis identifies damped quasi-periodic oscillations and allows estimation of the associated oscillatory parameters. We also observe an additional feature in the vortex residuals preceding the subsequent glitch, the origin of which remains unclear. These results motivate further theoretical and observational follow-up studies.
We explored several correlations and found a positive relationship between glitch magnitude and the subsequent inter-glitch interval, but only for large glitches. This suggests a possible connection between the magnitude of a glitch and the time until the next occurrence. However, this correlation does not hold when small glitches are taken into account. Consequently, a more comprehensive and robust study, with careful verification of small glitch events, is required. Furthermore, we estimated the pulsar’s braking index as 2.94 0.55.
@@
Future research includes investigating the post-glitch recovery and internal structure of other notable glitching pulsars, such as the Crab and PSR J1740–3015. To streamline the process, we are developing automated rotational parameter estimation scripts that perform the timing analysis quickly, enabling rapid model investigation and facilitating timely results.
@@ Acknowledgements
EG is supported by the Doctor Foundation of Qingdao Binhai University (No. BJZA2025025). BCJ acknowledges the support from Raja Ramanna Chair fellowship of the Department of Atomic Energy, Government of India (RRC – Track I Grant 3/3401 Atomic Energy Research 00 004 Research and Development 27 02 31 1002//2/2023/RRC/R&D-II/13886). BCJ also acknowledges support from the Department of Atomic Energy Government of India, under project number 12-R&D-TFR-5.02-0700. PA acknowledges the support from SERB-DST, Govt. of India, via project code CRG/2022/009359.
@@
Data and codes used for this research will be provided on reasonable request.
@@ Appendix A Results of Bayesian Analysis
Here, we present the outcomes of the Bayesian model comparison and the posteriors for the parameters estimated for each post-glitch recovery using the most preferred model. The results of the Bayesian Model comparison are listed in Tables 4 and 5. The posterior distributions are plotted with 68% and 95% credible intervals for all the cases given in Figs 14, 15, 16, 17, 18 and 19.
@@
| 57734 | 0.0 | 1283.0 | –3441.4 |
| Peak 1 Shifted | –0.1 | 886.1 | –3457.2 |
| 59417 | 0.0 | 3432.4 | –14222.3 |
| Peak 1 Shifted | –4.9 | 618.1 | –23907.7 |
| Peak 2 Shifted | –37.0 | 1581.4 | –14224.4 |